# A New Traders’ Game? — Response Functions in a Historical Perspective

### v1

@@ Abstract
Traders on financial markets generate non–Markovian effects in various ways, particularly through their competition with one another which can be interpreted as a game between different (types of) traders. To quantify the market mechanisms, we analyze self–response functions for pairs of different stocks and the corresponding trade sign correlators. While the non–Markovian dynamics in the self–responses is liquidity–driven, it is expectation–driven in the cross–responses which is related to the emergence of correlations. We study the non–stationarity of theses responses over time. In our previous analysis, we only investigated the crisis year 2008. We now considerably extend this by also analyzing the years 2007, 2014 and 2021. To improve statistics, we also work out averaged response functions for the different years. We find significant variations over time revealing changes in the traders’ game.
@@ 1 Introduction
There are two closely related empirical observations for individual assets that are essential to understand price formation. First, there are long–term memory effects in the order flow known as trade sign self–correlations or autocorrelations in the order flow [7, 8, 9, 10, 11, 12, 13]. They can be traced back to splitting of metaorders, also referred to as hidden orders. Second, there are self–responses that quantify the accumulated temporal influence of an incoming order and the subsequent price change of individual assets.
@@
These empirical findings can also be interpreted as a competition of two types of traders, liquidity takers and liquidity providers [1, 2, 3, 8, 9]. Liquidity takers split larger orders to not attract the attention by other traders with the intention to keep price levels around the midpoint constant. Their actions cause time–lagged correlations and thus trade sign self–correlations. Liquidity providers try to anticipate the liquidity takers’ trading behavior and take advantage of differences in the bid–ask spread by placing their orders at more expensive price levels. Liquidity takers are often willing to pay for the additional risk compensation. Due to non–stationarity in the traders‘ liquidity game the shape of self–responses are concave or increase for larger time lags [14]. However, the intraday price returns show a diffusive behavior if a certain balance between a sub– and superdiffusive dynamics exists for many stocks because traders exploit (statistical) arbitrage opportunities. Thus market efficiency is violated on shorter intraday time scales and restored on longer ones [8, 15].
@@
Similar empirical observations are also found when looking at different asset pairs for the same asset class. Empirical studies focus on the cross–response and the trade sign cross–correlations, respectively. Self– and cross–responses and trade sign correlators together dynamically reveal the non–Markovian effects that the market has on itself [16, 17, 18, 19, 20, 21, 22, 23]. These quantities are also closely related to the concept of cross–impact  [3, 19, 24, 25, 26, 27]. The cross–impact contains information on a significant fraction of return covariances [19]. Cross-response, trade sign cross–correlations and the cross–impact reflect the emergence of correlations which arise due to different traders’ market expectations and are thus expectation or information driven. The self–responses, trade sign self–correlations and self–impact are liquidity, i.e. supply and demand, driven [16, 17, 18, 19, 20, 21, 22, 23]. Correlations in the order flows of different stocks are caused by multi–asset trading such as portfolio rebalancing strategies and or pair trading [28]. The cross–impact is not directly needed for modeling correlations between returns and order flows of different stocks, but the cross–responses also contain important information on the traders’ actions, and thus on the non–stationarity in the trading.
@@
Here, we extend the analysis of Refs. [16, 17, 18] for the year 2008 by studying the years 2007, 2014 and 2021. There were two intriguing observations for the crisis year 2008. First, the whole market restored partially its efficiency in the course of a trading day, i.e. the cross–response for the market reached a maximum after some minutes, followed by a slowed decrease [16, 17]. Second, the memory in the order flow is long–term [18]. Here we show that the global crisis of 2008 had a major and lasting effect on self– and cross–responses. Our goal is to study this non–stationarity of the trading mechanism and to answer the question whether or not the traders’ game changed.
@@
After briefly introducing our data set in Sec. 2, we define self– and cross–responses and trade sign self– and cross–correlators in Sec. 3. The results of our study are presented in Sec. 4. We conclude with Sec. 5 by discussing in depth empirical evidence for a different traders’ behavior.
@@ 3 Response Functions and Trade Sign Correlators
If the price rises between two trades, the trading sign is $\varepsilon_{i}(t;n)=+1$. An increase in the price can result from a sellout of the entire volume at the best ask $a_{i}(t;n-1)$. Hence a trade sign $\varepsilon(t;n)=+1$ indicates that a trade was triggered by a buy order placed at time $(t;n-1)$. Similarly, a trade sign $\varepsilon(t;n)=-1$ means that a trade was triggered by a sell order at the time $(t;n-1)$. To obtain a trade sign for whole interval $t$, we sum up all $N(t)$ trade signs within a second
$\varepsilon_{i}(t)=\left\{\begin{array}[]{lll}\mathrm{Sgn}\left(\sum\limits_{n=1}^{N(t)}\varepsilon_{i}(t;n)\right)&\ ,&\mbox{if}\quad N(t)>0\\ 0&\ ,&\mbox{if}\quad N(t)=0\end{array}\,.\right.$ (5)
@@
$\mathrm{Sgn}(x)=\left\{\begin{array}[]{lll}+1&\ ,&\quad x>0\\ \hskip 7.68236pt0&\ ,&\quad x=0\\ -1&\ ,&\quad x<0\\ \end{array}\,.\right.$ (9)
@@ 4.5 Active and Passive Responses
$R_{i}^{(p)}(\tau)=\left\langle R_{ij}(\tau)\right\rangle_{j}\;,\quad\text{with$j\neq i$}$ (18)
@@
$R_{j}^{(a)}(\tau)=\left\langle R_{ij}(\tau)\right\rangle_{i}\;,\quad\text{with$i\neq j$}\,.$ (19)
The former is referred to as passive response function, the latter as active response function [18, 20]. The passive response function $R_{i}^{(p)}(\tau)$ quantifies the influence of all trade signs in the market on the return of stock $i$ with time lag $\tau$, the active response function $R_{j}^{(a)}(\tau)$ measures the influence of a trade or trades in stock $j$ on the returns of all other stocks with time lag $\tau$. In other words, index $i$ in $R_{i}^{(p)}(\tau)$ labels an influenced or impacted stock and index $j$ in $R_{j}^{(a)}(\tau)$ labels an influencing or impacting stock. Of course, every stock is influenced and is influencing at the same time.
@@
Figure 8: Passive and active cross-response functions $R_{i}^{(p)}(\tau)$ and $R_{j}^{(a)}(\tau)$ for $i,j=\textrm{AAPL},\textrm{GS},\textrm{XOM}$ in the years 2007 and 2008 versus time lag $\tau$. Negative values for passive and active cross–responses are removed.
@@
Figure 9: Passive and active cross-response functions $R_{i}^{(p)}(\tau)$ and $R_{j}^{(a)}(\tau)$ for $i,j=\textrm{AAPL},\textrm{GS},\textrm{XOM}$ in the years 2014 and 2021 versus time lag $\tau$. Negative values for passive and active cross–responses are removed.
@@ 5 Conclusions
We considerably extended the studies of Refs. [16, 17, 18] to investigate the non–stationarity of non–Markovian effects. We found in the dynamics of the trading sizeable changes in the dynamics which reveals severe modifications in the traders’ actions. The market self– and cross–responses and the market trade sign self– and cross–correlators facilitate, at least their averaged versions, rather general observations and statements. How do the different results fit together? To what extent is the crisis year 2008 special?
@@ Acknowledgment
We thank Shanshan Wang for fruitful discussions.
@@ Appendix A Industrial Sectors and Stocks
GD General Dynamics MTB M&\&&T Bank Corp.
@@
GENZ Genzyme Corp. FCX Freeport-McMoran Cp &\&& Gld
JNJ Johnson &\&& Johnson APD Air Products &\&& Chemicals
@@
CELG Celgene Corp. ROH Rohm &\&& Haas
@@
MA Mastercard Inc. PG Procter &\&& Gamble
@@
GD General Dynamics MTB M&\&&T Bank Corp.
@@
ABT Abbott Laboratories FCX Freeport-McMoran Cp &\&& Gld
JNJ Johnson &\&& Johnson APD Air Products &\&& Chemicals
@@
MA Mastercard Inc. PG Procter &\&& Gamble
@@
PCG PG&$\&E$ Corp CTL Century Telephone
@@
GD General Dynamics MTB M&\&&T Bank Corp.
@@
BIIB Biogen Idec Inc.. FCX Freeport-McMoran Cp &\&& Gld
JNJ Johnson &\&& Johnson APD Air Products &\&& Chemicals
@@
MA Mastercard Inc. PG Procter &\&& Gamble
@@
GD General Dynamics MTB M&\&&T Bank Corp.
@@
BIIB Biogen Idec Inc.. FCX Freeport-McMoran Cp &\&& Gld
JNJ Johnson &\&& Johnson APD Air Products &\&& Chemicals
@@
MA Mastercard Inc. PG Procter &\&& Gamble

### v2

# A New Traders’ Game? — Empirical Analysis of Response Functions in a Historical Perspective

@@ Abstract
Traders on financial markets generate non–Markovian effects in various ways, particularly through their competition with one another which can be interpreted as a game between different (types of) traders. To quantify the market mechanisms, we empirically analyze self–response functions for pairs of different stocks and the corresponding trade sign correlators. While the non–Markovian dynamics in the self–responses is liquidity–driven, it is expectation–driven in the cross–responses which is related to the emergence of correlations. We empirically study the non–stationarity of these responses over time. In our previous data analysis, we only investigated the crisis year 2008. We now considerably extend this by also analyzing the years 2007, 2014 and 2021. To improve statistics, we also work out averaged response functions for the different years. We find significant variations over time revealing changes in the traders’ game.
@@ 1 Introduction
In the research field of market microstructure [7, 8, 9], one of the first models of how trading mechanisms influence price formation was put forward by Kyle [10]. In his foundational paper, the author models the interplay of stylized types of traders, namely noise traders, a trader with insider information and a market maker. From this he deduces that the price impact is scaling linear with the order flow. Glosten and Milgrom model the influence of trades on the price formation from the viewpoint of market makers, who want to protect themselves from better informed traders [11]. Their work connected the economic terms of adverse selection and information asymmetry with the context of market microstructure. Some more recent modeling efforts can be found in Refs. [12, 13]. In the former, the authors use an analytically tractable Markov queueing process to model a stylized order flow. The latter models price changes and order flow with a multivariate Hawkes process, to capture the self–exciting nature of financial markets.
@@
There are two closely related empirical observations for individual assets that are essential to understand price formation. First, there are long–term memory effects in the order flow known as trade sign self–correlations or autocorrelations in the order flow [14, 15, 16, 17, 18, 19]. They can be traced back to splitting of metaorders, also referred to as hidden orders. Second, there are self–responses that quantify the accumulated temporal influence of an incoming order and the subsequent price change of individual assets. In Refs. [20, 21] the authors show empirical evidence for both, strong autocorrelations between daily imbalances of buy and sell orders, as well as correlations between returns and lagged order imbalance on a horizon of one trading day. In Ref. [22] the authors introduce the order flow imbalance, an empirical variable which entails changes in bid and ask queues, and provide robust evidence of a linear relation between this variable and returns on a high–frequency timescale.
@@
These empirical findings can also be interpreted as a competition of two types of traders, liquidity takers and liquidity providers [1, 2, 3, 14]. Liquidity takers split larger orders to not attract the attention by other traders with the intention to keep price levels around the midpoint constant. Their actions cause time–lagged correlations and thus trade sign self–correlations. Liquidity providers try to anticipate the liquidity takers’ trading behavior and take advantage of differences in the bid–ask spread by placing their orders at more expensive price levels. Liquidity takers are often willing to pay for the additional risk compensation. Due to non–stationarity in the traders‘ liquidity game the shape of self–responses are concave or increase for larger time lags [23]. However, the intraday price returns show a diffusive behavior if a certain balance between sub– and superdiffusive dynamics exists for many stocks because traders exploit (statistical) arbitrage opportunities. This was analyzed in Ref. [24], where the authors introduce the virtual price impact, the impact a trade would have solely due to the current state of the order book. They show that the volume imbalance of limit buy vs sell orders is negatively correlated to the returns. This explains the difference between real and virtual price impact, as incoming limit orders cancel the predictable change in price due to trades. Thus market efficiency is violated on shorter intraday time scales and restored on longer ones [14, 25].
@@
Similar empirical observations are also found when looking at different asset pairs for the same asset class. In the economics literature various kinds of cross–effects were identified in Refs. [26, 27, 28, 29]. In the econophysics literature, empirical studies focus on the cross–response and the trade sign cross–correlations, respectively. Self– and cross–responses and trade sign correlators together dynamically reveal the non–Markovian effects that the market has on itself [30, 31, 32, 33, 34, 35, 36, 37]. These quantities are also closely related to the concept of cross–impact  [3, 33, 27, 38, 39, 40]. The cross–impact contains information on a significant fraction of return covariances [33]. Cross-response, trade sign cross–correlations and the cross–impact reflect the emergence of correlations which arise due to different traders’ market expectations and are thus expectation or information driven. The self–responses, trade sign self–correlations and self–impact are liquidity, i.e. supply and demand, driven [30, 31, 32, 33, 34, 35, 36, 37]. Correlations in the order flows of different stocks are caused by multi–asset trading such as portfolio rebalancing strategies and or pair trading [41]. The cross–impact is not directly needed for modeling correlations between returns and order flows of different stocks, but the cross–responses also contain important information on the traders’ actions, and thus on the non–stationarity in the trading. Further details have been explored, such as the market collective response [42] and the role of asymmetric information in market impacts [43]
@@
Here, we extend the analysis of Refs. [30, 31, 32] for the year 2008 by studying the years 2007, 2014 and 2021. In doing so, we rely on the well–established and theoretically explored framework of self– and cross–response functions in returns and trade signs [30, 31, 32, 33, 34, 36, 37], which we apply in a large scale empirical analysis across the several distinct time periods. The comparison of the years mentioned above covers pre–crisis, crisis and post–crisis periods and also includes the rise of algorithmic trading. We empirically analyze and discuss our results regarding both, the market, as well as the sectoral level, and distinguish between intra–sectoral and inter–sectoral measures. There were two intriguing observations for the crisis year 2008. First, the whole market restored partially its efficiency in the course of a trading day, i.e. the cross–response for the market reached a maximum after some minutes, followed by a slowed decrease [30, 31]. Second, the memory in the order flow is long–term [32]. Here we show that the global crisis of 2008 had a major and lasting effect on self– and cross–responses. Our goal is to provide an in depth empirical account using established empirical measures of market microstructure to study this non–stationarity of the trading mechanism over different periods and to answer the question whether or not the traders’ game changed.
@@
After briefly introducing our data set in Sec. 2, we define self– and cross–responses and trade sign self– and cross–correlators in Sec. 3. The results of our data analysis are presented in Sec. 4. Theoretical and Modeling Aspects are briefly discussed in Sec. 5. We conclude with Sec. 6 by discussing in depth empirical evidence for a different traders’ behavior.
@@ 3 Response Functions and Trade Sign Correlators
If the price rises between two trades, the trading sign is $\varepsilon_{i}(t;n)=+1$. An increase in the price can result from a sellout of the entire volume at the best ask $a_{i}(t;n-1)$. Hence a trade sign $\varepsilon(t;n)=+1$ indicates that a trade was triggered by a buy order placed at time (t;n1)(t;n-1). Similarly, a trade sign $\varepsilon(t;n)=-1$ means that a trade was triggered by a sell order at the time (t;n1)(t;n-1). To obtain a trade sign for whole interval $t$, we sum up all $N(t)$ trade signs within a second
$\varepsilon_{i}(t)=\left\{\begin{array}[]{lll}\mathrm{Sgn}\left(\sum\limits_{n=1}^{N(t)}\varepsilon_{i}(t;n)\right)&\ ,\quad&\mbox{if}\quad N(t)>0\\ 0&\ ,\quad&\mbox{if}\quad N(t)=0\end{array}\,.\right$. (5)
@@
$\mathrm{Sgn}(x)=\left\{\begin{array}[]{lll}+1&\ ,\quad&\quad x>0\\ \hskip 7.68236pt0&\ ,\quad&\quad x=0\\ -1&\ ,\quad&\quad x<0\\ \end{array}\,.\right$. (9)
@@ 4.5 Active and Passive Responses
$\overline{R}_{i}^{(p)}(\tau)=\left\langle R_{ij}(\tau)\right\rangle_{j}\;,\quad\text{with $j\neq i$}$ (18)
@@
$\overline{R}_{j}^{(a)}(\tau)=\left\langle R_{ij}(\tau)\right\rangle_{i}\;,\quad\text{with $i\neq j$}\,.$ (19)
The former is referred to as passive response function, the latter as active response function [32, 34]. The passive response function $\overline{R}_{i}^{(p)}(\tau)$ quantifies the influence of all trade signs in the market on the return of stock $i$ with time lag $\tau$, the active response function $\overline{R}_{j}^{(a)}(\tau)$ measures the influence of a trade or trades in stock $j$ on the returns of all other stocks with time lag $\tau$. In other words, index $i$ in $\overline{R}_{i}^{(p)}(\tau)$ labels an influenced or impacted stock and index $j$ in $\overline{R}_{j}^{(a)}(\tau)$ labels an influencing or impacting stock. Of course, every stock is influenced and is influencing at the same time.
@@
Figure 8: Passive and active cross-response functions $\overline{R}_{i}^{(p)}(\tau)$ and $\overline{R}_{j}^{(a)}(\tau)$ for $i,j=\textrm{AAPL},\textrm{GS},\textrm{XOM}$ in the years 2007 and 2008 versus time lag $\tau$. Negative values for passive and active cross–responses are removed.
@@
Figure 9: Passive and active cross-response functions $\overline{R}_{i}^{(p)}(\tau)$ and $\overline{R}_{j}^{(a)}(\tau)$ for $i,j=\textrm{AAPL},\textrm{GS},\textrm{XOM}$ in the years 2014 and 2021 versus time lag $\tau$. Negative values for passive and active cross–responses are removed.
@@
5 Theoretical and Modeling Aspects
In the pioneering works on response functions [14] and price impacts [15], two at first sight competing, yet largely consistent, explanations for the non–Markovian effects were offered. In Ref. [14], the self–responses (response of the price to trading of this stock) were introduced and successfully used to uncover non–Markovian effects on intraday timescales. In Ref. [15], long–term memory effects in order flow were analyzed by studying autocorrelations of order signs. In line with this, the authors investigated how liquidity fluctuations function to maintain market efficiency.
Both explanations were formulated in mathematical models. Ref. [14] introduces the transient price impact model and provides explicit formulae for the self–response whose key ingredient is a “bare” self–impact function depending on the time lag $\tau$. The shape of this function encodes how the non–Markovian effects act, in Ref. [14] an algebraic form is assumed, corresponding to a powerlaw dependence on the time lag. We mention in passing and to avoid confusion in the terminology that this bare self–impact function formally plays a role comparable to the “response function” in response theory as applied in traditional statistical mechanics and condensed matter theory. Here, we use the term response function in a related, but different way as defined in the cited econophysics literature. The qualitative introduction and discussion of the permanent price impact model in Ref. [15] was complemented by a mathematical model compared to the transient price impact model [1]. A related model for order splitting was already put forward in Ref. [16].
All these studies were confined to the self–responses or self–impact, respectively. The early studies in the economics literature on various cross–effects [26, 27, 28, 29] also contain modeling aspects. The first empirical results on the cross–response functions and the trade–sign cross–correlations [30, 31, 32] prompted theoretical work and extensions of the above models. Simultaneously but independently, such studies were put forward in Refs. [33] and [34]. For the cross–response functions they involve a “bare” cross–impact function depending on the time lag $\tau$. While the bare self–impact is driven by non–Markovian liquidity effects, the bare cross–impact function must be entirely information–driven. Detailed empirical analyses were also carried out in Refs. [33] and [34]. Various details in the fine structure of cross–responses and trade–sign cross–correlations were subsequently investigated. In Ref. [42], the market collective responses were statistically analyzed employing singular value decomposition. The crucial role of the traded volume was revealed which in turn is related to liquidity changes and imbalances. In Ref. [43], asymmetry in the information affecting market impacts was analyzed and the model of Ref. [34] was extended accordingly.
We also mention briefly that there has been substantial effort in extending theoretical frameworks from thermodynamics and dynamical systems to incorporate non–equilibrium dissipation and non–stationarity [50, 51]. This has led, for example, to generalized fluctuation relations and the introduction of dissipation functions as theoretical tools to describe the evolution of systems in traditional physics. While these developments provide valuable conceptual background, our approach here does not build on such general formulations. This would be rewarding but a formidable task, as the fundamental equations describing the dynamics of financial markets are not known, and the underlying mechanisms themselves change over time. Instead, we focus on data analysis and the established response function methodology of econophysics. In a statistical setting for complex systems, the approach of [34] provides information on the dynamical evolution of traders’ behavior and their interactions in a single stock as well as across different stocks. Both short–term price changes, resulting from temporary liquidity shortages, and long–term price changes, which may be driven by private information, significantly influence trading behavior.
@@ 6 Conclusions
We considerably extended the studies of Refs. [30, 31, 32] to investigate the non–stationarity of non–Markovian effects. We found in the dynamics of the trading sizable changes in the dynamics which reveals severe modifications in the traders’ actions. The market self– and cross–responses and the market trade sign self– and cross–correlators facilitate, at least their averaged versions, rather general observations and statements. How do the different results fit together? To what extent is the crisis year 2008 special?
@@ Acknowledgment
We thank Shanshan Wang for fruitful discussions, we are particularly grateful to Rüdiger Kiesel for helpful remarks on the economics literature.
@@ Appendix A Industrial Sectors and Stocks
GD General Dynamics MTB M&\&T Bank Corp.
@@
GENZ Genzyme Corp. FCX Freeport-McMoran Cp &\& Gld
JNJ Johnson &\& Johnson APD Air Products &\& Chemicals
@@
CELG Celgene Corp. ROH Rohm &\& Haas
@@
MA Mastercard Inc. PG Procter &\& Gamble
@@
GD General Dynamics MTB M&\&T Bank Corp.
@@
ABT Abbott Laboratories FCX Freeport-McMoran Cp &\& Gld
JNJ Johnson &\& Johnson APD Air Products &\& Chemicals
@@
MA Mastercard Inc. PG Procter &\& Gamble
@@
PCG PG&E\&E Corp CTL Century Telephone
@@
GD General Dynamics MTB M&\&T Bank Corp.
@@
BIIB Biogen Idec Inc.. FCX Freeport-McMoran Cp &\& Gld
JNJ Johnson &\& Johnson APD Air Products &\& Chemicals
@@
MA Mastercard Inc. PG Procter &\& Gamble
@@
GD General Dynamics MTB M&\&T Bank Corp.
@@
BIIB Biogen Idec Inc.. FCX Freeport-McMoran Cp &\& Gld
JNJ Johnson &\& Johnson APD Air Products &\& Chemicals
@@
MA Mastercard Inc. PG Procter &\& Gamble